Subject Summary

If you missed some lectures, then don’t miss this

one!

             Nick Urbanik <nicku@nicku.org>
Copyright Conditions: GNU FDL (see http://www.gnu.org/licenses/fdl.html  )

A computing department

Main Topics in this Subject

Perl snmp Network Troubleshooting Network Design

dhcp and dns systems

ldap directories and querying them Routing, Route Summarisation

Perl

We covered the basics of the language Refer to the Perl Summary that I wrote You did some laboratory exercises on Perl

Revise your work carefully in all of them Detailed solutions are on the website

Make sure you know how to write a subroutine and how to call it, and how to pass parameters I will provide a copy of the booklet Perl Reference Quide in the exam (no need to bring your own copy)

Please return the booklet in the exam Please do not write in it or mark it

There is one question in the exam on this topic

SNMP

There are two sets of notes on this topic; one (short set) from Albert on the iso network management model

don’t just memorise the names; sorry, you need to understand what you study!

one (long set) on the practicalities of snmp; read them carefully

You put some effort and time into Cricket in the lab. Make sure you know what you were doing, and what Cricket can do.

Make sure that you understand all the details of the snmp operations (requests); read the labs again.

There is one question on this topic in the exam

Network Troubleshooting — 1

This is a new topic; there are no previouos exam questions on this topic.

I focussed on the use of Ethereal and tcpdump for collecting information about network problems

However, Ethereal screen dumps don’t fit well into an exam paper I wrote additional notes on using tcpdump with dhcp

Some of these may be included as an appendix in the exam; understand, don’t memorise.

So you may use your judgement as to which Ethernet packet sniffing tool I refer to in the exam.

Network Troubleshooting — 2

We also examined many uses of such important and commonly available basic tools as traceroute (or

tracert) and ping

Do not waste time memorising any lengthy formulas; if they are required, they will appear in an appendix to the exam.

We covered many topics in this area, including troubleshooting a switched network

Altogether, there is one question directly on this area, and half a question that is at least partly on this area

dhcp and dns Management Systems

We covered the basics of dhcp operation including normal operation, and problems that can occur in dhcp operation

We also studied two schemes for automating dhcp and dns management

One scheme introduced at this Institute, in the Computer Centre and our Department, One fully automated registration scheme involving a web registration application to which new users are automatically directed

There is half a question directly on these topics, and another half question that includes these topics

ldap directories and querying them — 1

We covered a number of aspects of ldap directories, and

I provided a laboratory exercise in setting up network accounts on a directory server, But due to insufficient laboratory time, we did not work through this in the classroom That will not be in the exam

However, we covered a few topics in authentication and authorisation

both in the lectures and in the laboratory

ldap directories and querying them — 2

I have already told you that:

an appendix to the exam paper from rfc 2254 describes the grammar for ldap Another appendix to the exam paper from rfc 2255 describes the grammar for ldap urls.

There is one question on this topic in the exam.

Routing, Route Summarisation

Due to the sars epidemic, my planned coverage of this area has been reduced, mainly to covering:

route summarisation and

address block allocation and design.

In particular, we have not covered the same areas of switching as in the previous year

I suggest that you be guided by the tutorial exercises we do in this area.

There is half a question on this topic in the exam.

Format of the Exam (2002–2003)

There are five questions There are no options—do all five questions That’s good—no need to waste time deciding which one to leave out!

Advice for the Exam

Budget your time wisely in the exam: Spend a few minutes to plan your time Divide remaining time by five Do not spend more than this time on any one question until you have answered five questions fully

Show your working A wrong answer with no working gets zero marks A wrong answer with some working that is on the right track gets some marks

Summary of Topics in Exam

Topic Number of Questions: directly, partly

Perl  1
snmp  1
Network Troubleshooting  1, ^{1}/_{2}

Network Design:

dhcp and dns systems  ^{1}/_{2}, ^{1}/_{2}
ldap directories and querying them  1
Routing, Route Summarisation  ^{1}/_{2}

Watch the Subject Web Site

I will begin writing solutions to past exam papers, the test, laboratory exercises and workshops, and post them on our subject web site

Watch the web site for announcements:

I will write and post solutions to problems as soon as I can.

I will make a new icon to highlight changes on the site, including solutions to problems as I write them.